Assume a population of 1,3, and 8. Assume that samples of size n=2 are randomly selected with replacement from the population.

Sample   x1   x2
1   1   1
2   1   3
3   1   8
4   3   1
5   3   3
6   3   8
7   8   1
8   8   3
9   8   8

Find the value of the population standard deviation sigmaσ.
Answer #1

Question asked:

Value of the population standard deviation \sigma:

Answer:

Population values: 1, 3, 8

From the given population data, the following Table is calculated:

Population Mean (\mu) is given by:

\mu =\frac{\sum_{i=1}^{3}x_{i}}{3}=\frac{12}{3}=4

x x - \mu (x - \mu)2
1 - 3 9
3 - 1 1
8 4 16
Total = 26

Population Standard Deviation (\sigma) is given by:

Σǐel (ri-μ)2 /26 = V8.6667= 2.9439

So,

Answer is:

2.9439